Write don the four maxwell's equations in a region of space with charge density d and current density J and discuss qualitatively what each equation describing
1
Expert's answer
2020-08-21T10:57:45-0400

J = Time varying current

H = Time varying magnetic field

E = electric filed

D = Electric flux densities (volt/m)

B =Magnetic flux densities (ampere/m)


Maxwell’s equations for all electromagnetic phenomena:


Maxwell's equation between E and B (Faraday’s law of induction)

"\triangledown\times E = - \frac{\partial B}{\partial t}" ...1)


Maxwell's equation between H ,J and D(Amp`ere’s law )

"\triangledown\times H = J + \frac{\partial D}{\partial t}" ...2)


Maxwell's equation between D and "\rho"


"\triangledown. D = \rho" ...3) (Gauss’ laws)


"\triangledown. B = 0" ... 4) (Gauss’ laws)


The first is Faraday’s law of induction, the second is Amp`ere’s law as amended by Maxwell to include the displacement current ∂D/∂t, the third and fourth are Gauss’ laws for the electric and magnetic fields.